OpenStudy (anonymous):

Find a particular solution of y'' - y = e^x by the method of undetermined coefficients using the guess yp = Axe^x.

OpenStudy (jamesj):

So substitute and evaluate ... y' = A(x+1)e^x y'' = A(x+2)e^x Hence y'' - y = ... = e^x

OpenStudy (jamesj):

=> A = ...